OpenStudy (anonymous):

yolanda invested $2,000 in a us treasury note that pays 5% simple interest per year . how much interest did yolanda earn after 3 years is she does not cash in her notes??/

OpenStudy (anonymous):

A = P + Prt = 2000 + 2000(0.05)(3)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

what!!??

OpenStudy (jdoe0001):

i = prt

OpenStudy (ranga):

They are asking for the simple interest earned: Simple interest = P * r * t ----- (1) P = $2,000 r = 5% = 0.05 t = 3 years Substitute the numbers into (1)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so the answer 100

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

1200

OpenStudy (ranga):

2000 x 0.05 x 3 = ?

OpenStudy (anonymous):

300

OpenStudy (ranga):

Yes. $300.